The length of a rectangle is 1 in. more than twice its width. If the perimeter of the rectangle is 32 in., find the width of the rectangle.
Answer by checkley71(8403)

LENGTH=2W+1
WIDTH=2L+2W=32
2(2W+1)+2W=32
4W+2+2W=32
6W=32-2
6W=30
W=30/6
W=5 IS THE WIDTH.
L=2*5+1
L=10+1
L=11 IS THE LENGTH.
PROOF
2*5+2*11=32
10+22=32
32=32